# Steganography Beyond Space-Time With Chain of Multimodal AI Agents

### v1

@@ Abstract
Steganography is the art and science of covert writing, with a broad range of applications interwoven within the realm of cybersecurity. As artificial intelligence continues to evolve, its ability to synthesise realistic content emerges as a threat in the hands of cybercriminals who seek to manipulate and misrepresent the truth. Such synthetic content introduces a non-trivial risk of overwriting the subtle changes made for the purpose of steganography. When the signals in both the spatial and temporal domains are vulnerable to unforeseen overwriting, it calls for reflection on what can remain invariant after all. This study proposes a paradigm in steganography for audiovisual media, where messages are concealed beyond both spatial and temporal domains. A chain of multimodal agents is developed to deconstruct audiovisual content into a cover text, embed a message within the linguistic domain, and then reconstruct the audiovisual content through synchronising both aural and visual modalities with the resultant stego text. The message is encoded by biasing the word sampling process of a language generation model and decoded by analysing the probability distribution of word choices. The accuracy of message transmission is evaluated under both zero-bit and multi-bit capacity settings. Fidelity is assessed through both biometric and semantic similarities, capturing the identities of the recorded face and voice, as well as the core ideas conveyed through the media. Secrecy is examined through statistical comparisons between cover and stego texts. Robustness is tested across various scenarios, including audiovisual compression, face-swapping, voice-cloning and their combinations.
Index Terms:
Artificial intelligence, deepfakes, multimodality, steganography.
@@ I Introduction
Steganography is the study of covert writing, which has evolved from rudimentary arts, such as the use of invisible ink, into a sophisticated scientific discipline, interwoven within the field of cybersecurity [1, 2, 3]. The applications of steganography are vast and varied, encompassing secret communication [4, 5, 6, 7, 8, 9], anti-counterfeit watermarking [10, 11, 12, 13, 14], provenance tracking [15, 16, 17] and forensic analysis [18, 19, 20], among others [21, 22, 23]. In general, a steganographic system involves the process of embedding a message into a cover medium and then extracting the hidden message from the resulting stego medium. Steganography manifests in various forms, including visual, aural and linguistic content. In other words, information can be concealed in various aspects of digital media, such as pixel intensities in imagery [24], sound waves in audio [25], synonym words in text [26] and frequency coefficients for transformed media [27]. The distortion introduced by steganographic systems is typically constrained to remain imperceptible to human senses. The fundamental objective of steganography is to keep the hidden information invisible and inaudible to all except the authorised parties who possess the key.
@@
Consider a human-centric audiovisual content in which an individual’s speech was recorded both aurally and visually. Suppose that the message is hidden within the pixel intensities or sound waves, serving an arbitrary purpose. Although steganographic systems can be designed to withstand distortion caused by common signal processing operations such as compression, there remains a risk of unforeseen manipulations that go beyond the robustness premise. Generative AI technologies, such as face-swapping and voice-cloning, can potentially wipe out the hidden message, as the synthetic content has a non-trivial chance of overwriting the subtle changes made for the purpose of steganography.
@@
If the signals in both spatial and temporal domains are at risk of unforeseen manipulations, it prompts a reconsideration of what remains invariant after all. In quest for an invariant domain, this study proposes a paradigm in steganography where messages are concealed beyond spatial and temporal domains. Motivated by multi-agent collaboration [61], we develop a chain of multimodal AI agents to deconstruct audiovisual content into a cover text, encode a message within the linguistic domain, and subsequently reconstruct the audiovisual content by synchronising the visual and aural modalities with the resulting stego text. The foundation of applied linguistic steganographic coding involves biasing the statistical distribution of word sampling [62]. Suppose a set of keywords (tokens) is shared between the sender and the receiver as the stego key. The message is encoded by biasing the word sampling process of a language generation model tasked with paraphrasing the given text, steering the model towards generating words from the shared set. The message can later be decoded by analysing the statistical likelihood of word choices, in comparison to the expected distribution of the shared set. The accuracy of message transmission is evaluated in both zero-bit and multi-bit capacity settings. In contrast to the perceptual similarity typically adopted in prior work, fidelity is extended to encompass both biometric and semantic similarities, representing identities of the recorded face and voice, as well as the ideas conveyed through the media. Secrecy is assessed through statistical analysis of cover and stego texts. Robustness is tested under conditions such as audiovisual compression, face-swapping, voice-cloning, and combined scenarios.
@@
The remainder of this paper is organised as follows. Section II outlines the scope of this study in relation to the primary properties of steganography. Section III presents the proposed methodology, which involves a chain of multimodal AI agents. Section IV details the experimental implementation and discusses the evaluation results. Finally, Section V concludes the paper with a summary of the research findings and potential directions for future research.
II Preliminaries
@@ II-A Capacity
II-B Fidelity
Fidelity refers to the degree of similarity between the cover and stego versions of the medium. A typical fidelity requirement is based on perceptual similarity, guaranteeing that any distortion caused by the steganographic system is either invisible or inaudible to human perception. The steganographic system should not compromise the integrity of the auditory or visual quality to the point of being noticeable to the listener or viewer. In a broader sense, however, the concept of fidelity can be relaxed to refer to semantic similarity, where the alterations made to the medium may be noticeable but do not change the fundamental meaning or purpose conveyed. This is often exemplified in linguistic steganography, where changes in lexicon, syntax, or even language may result in synonymous expressions or paraphrases that, while recognisable to a human observer, retain the core idea and intent of the original text. For human-centric audiovisual media, fidelity can be broadened to encompass biometric similarity, including face identity for visual content and voice identity for aural content. This study focuses on both semantic and biometric similarities in the context of multimodal steganography.
II-C Secrecy
@@
II-D Robustness
Robustness refers to the ability to reliably decode the message even in the presence of unintentional degradations or malicious operations. It is pivotal in applications such as fact-checking and cyber-forensics, as it enables the tracing of a content’s origin and provides a means to verify whether the media has been fabricated or misrepresented. Central to achieving robustness is the principle of orthogonality, which ensures that the embedded message remains unaffected and detectable despite a range of potential manipulations. When the message is embedded in a domain that is either orthogonal or invariant to the potential attacks, it remains immune to their effects. For example, a visual or aural medium can be projected into the frequency domain, where the message is embedded in parts of the frequency spectrum that are less likely to be compromised by potential manipulations. This study exploits the orthogonality between the linguistic and audiovisual domains, allowing the message to withstand changes in either the aural or visual modality.
@@ III Methodology
III-A Message Encoding Process
The encoding process begins with demultiplexing the cover multimedia container into cover video and cover audio. Next, the cover audio is transcribed into cover text, which is then encoded into stego text using the shared key. In the zero-bit capacity setting, only one key $K$ is shared. In the multi-bit capacity setting, the key $K$ to be used is selected from a set of shared keys based on the corresponding message symbol $S$, where the size of the message alphabet $\|\mathcal{S}\|$ matches the size of the key set $\|\mathcal{K}\|$. The stego text is then narrated into stego audio, and the cover video is synchronised with the stego audio. Finally, both stego video and stego audio are multiplexed into a stego multimedia container. An overview of the message encoding process is illustrated in Figure 1.
@@ III-A Message Encoding Process
The cover audio stream $A$ is transcribed into the cover text transcript $T$ using a speech-to-text (STT) agent:
@@
Consider a language generation agent prompted to paraphrase the given text. The word sampling process during paraphrasing is parameterised by a shared key $K$, selected based on the intended message symbol. The language generation agent paraphrases the cover text transcript $T$ and results in a stego text transcript $T^{\prime}$:
@@
The word sampling process is biased toward generating tokens in the keyword set $\mathcal{W}$. The output of a language generation agent is a sequence of logits, representing unnormalised probabilities of selecting each token from the dictionary. Let $z_{w}$ denote the logit for the token $w$. The probability of selection is biased by adjusting the logit:
@@
$p_{w}=\frac{\exp(z^{\prime}_{w})}{\sum_{w\in\mathcal{V}}\exp(z^{\prime}_{w})}.$ (6)
As long as the size of the keyword set is large enough, it is possible to preserve semantic equivalence between the cover and stego transcripts, with a statistically significant portion of words sampled from the keyword set.
@@
The stego text transcript $T^{\prime}$ is converted into the aural modality using the voice cloned from the cover audio stream $A$ through a text-to-speech (TTS) agent:
@@
To maintain audiovisual consistency, the stego audio stream $A^{\prime}$ is synchronised with the cover video stream $V$ using a lip-synchronisation agent:
@@
III-B Message Decoding Process
@@
The query audio stream $\hat{A}$ is transcribed into a query text transcript $\hat{T}$ using an STT agent:
@@
IV Evaluations
@@ IV-A Experimental Setup
  • Language Generator: An open-source language generation model named Llama developed by Meta.
These models were integrated to enable a fully automated steganographic system capable of generating semantics-aligned identity-preserved audiovisual media. The evaluations on accuracy, fidelity and robustness were conducted on a 9-second speech by C. E. Shannon from the MIT centennial film ‘The Thinking Machine’ (1961), which served as the cover medium. The evaluations on capacity and secrecy were conducted on 100 paragraphs of approximately 50 words each from ‘Alice’s Adventures in Wonderland’ by Lewis Carroll.
Figure 3: Evaluation of accuracy in zero-bit capacity setting.
Figure 4: Evaluation of accuracy in multi-bit capacity setting.
IV-B Evaluation of Zero-Bit Capacity
@@
IV-C Evaluation of Multi-Bit Capacity
@@
IV-D Evaluation of Fidelity
@@
Figure 7: Evaluation of secrecy with respect to perplexity.
IV-E Evaluation of Secrecy
At first glance, the challenge of steganalysis may appear to be a problem of generative AI detection. However, AI-generated content does not necessarily imply concealed communications, and it is impractical to classify all AI-generated content as stego content. In contrast, we evaluate the statistical deviations between cover and stego texts. Figure 6 shows the frequency distribution of words in the cover and stego texts, plotted on a log-log scale. The distribution for both cover and stego texts follows a similar pattern, with a steep decline in frequency as rank increases, typical of a Zipfian distribution [69]. Figure 7 presents the perplexity distribution for both cover and stego texts, measured using the Llama language model. Both cover and stego texts exhibit a similar range and central tendency in perplexity values, although the stego texts contain more outliers. These findings suggest that, providing the stego texts are not outliers, they can be sampled with the same level of uncertainty as the cover texts, as reflected in their perplexity distribution.
@@ IV-E Evaluation of Secrecy
Figure 9: Visualisation of cover, stego, compression and deepfake versions of media across visual, aural and linguistic modalities.
IV-F Evaluation of Robustness
Figure 8 presents the robustness of the steganographic system under different conditions: no manipulation, compression, deepfake, and hybrid (both deepfake and compression) scenarios. The compression algorithm reduced the display resolution to 1/16, from 320 × 240 to 80×60 pixels, halved the video frame rate from 14.92 to 7.46 frames per second (fps), and halved the audio sample rate from 44,100 Hz to 22,050 Hz. The deepfake video streams were generated using the face-swapping model SimSwap [70], with 100 faces from the FaceForensics++ dataset [71], while the deepfake audio streams were generated using the voice-cloning model XTTS [64], with 100 voices from the LibriSpeech corpus [72]. In the compression scenario, the TPR was 0.95, the same as in the no-manipulation scenario. In the deepfake scenario, the TPR was 0.94, showing only a negligible decrease compared to the no-manipulation scenario, likely due to the unclear speech generated by the voice-cloning model and the imperfections in the speech-to-text model. Overall, these results demonstrate the system’s high robustness across various forms of media manipulation. Figure 9 visualises a pair of cover and stego media, along with the compression and deepfake variants. It can be observed that while the cover transcript was altered by the steganographic system, the core semantics were preserved in the stego transcript. For instance, the text changed from ‘I confidently expect that within a matter of 10 or 15 years, something will emerge from the laboratories, which is not too far from the robot of science fiction fame’ to ‘Scientists are likely to create a new artificial life form that closely resembles the intelligent machines of science fiction’.
V Conclusion
In conclusion, this study investigates the vulnerability of steganography to the evolving capabilities of generative AI, with a focus on the risk of synthetic content overwriting hidden messages. Instead of relying on the spatial or temporal domains, messages are embedded in the linguistic domain of audiovisual content based on a chain of multimodal AI agents. At the core of this chain is a language generation model tasked with paraphrasing the given transcript, while the word sampling process is biased towards a shared set of tokens. The paraphrased transcript is then converted into sound using a voice-cloning model, and the video is aligned with the audio using a lip-synchronisation model. Several aspects of the proposed steganographic system are evaluated, including accuracy, fidelity, secrecy and robustness, across various metrics and conditions. Future research may explore the integration of steganographic methods for individual modalities into a unified framework, where they complement each other through the cooperation and interaction of multimodal AI agents.

### v2

@@ Abstract
Steganography is the art and science of covert writing, with a broad range of applications interwoven within the realm of cybersecurity. As artificial intelligence continues to evolve, its ability to synthesise realistic content emerges as a threat in the hands of cybercriminals who seek to manipulate and misrepresent the truth. Such synthetic content introduces a non-trivial risk of overwriting the subtle changes made for the purpose of steganography. When the signals in both the spatial and temporal domains are vulnerable to unforeseen overwriting, it calls for reflection on what, if any, remains invariant. This study proposes a paradigm in steganography for audiovisual media, where messages are concealed beyond both spatial and temporal domains. A chain of multimodal artificial intelligence is developed to deconstruct audiovisual content into a cover text, embed a message within the linguistic domain, and then reconstruct the audiovisual content through synchronising both auditory and visual modalities with the resultant stego text. The message is encoded by biasing the word sampling process of a language generation model and decoded by analysing the probability distribution of word choices. The accuracy of message transmission is evaluated under both zero-bit and multi-bit capacity settings. Fidelity is assessed through both biometric and semantic similarities, capturing the identities of the recorded face and voice, as well as the core ideas conveyed through the media. Secrecy is examined through statistical comparisons between cover and stego texts. Robustness is tested across various scenarios, including audiovisual resampling, face-swapping, voice-cloning and their combinations.
@@ I Introduction
Steganography is the study of covert writing, which has evolved from rudimentary arts, such as the use of invisible ink, into a sophisticated scientific discipline, interwoven within the field of cybersecurity [1, 2, 3]. The applications of steganography are vast and varied, encompassing secret communication [4, 5, 6, 7, 8, 9], anti-counterfeit watermarking [10, 11, 12, 13, 14], provenance tracking [15, 16, 17] and forensic analysis [18, 19, 20], among others [21, 22, 23]. In general, a steganographic system involves the process of embedding a message into a cover medium and then extracting the hidden message from the resulting stego medium. Steganography manifests in various forms, including visual, auditory and linguistic content. In other words, information can be concealed in various aspects of digital media, such as pixel intensities in imagery [24], sound waves in audio [25], synonym words in text [26] and frequency coefficients for transformed media [27]. The distortion introduced by steganographic systems is typically constrained to remain imperceptible to human senses. The fundamental objective of steganography is to keep the hidden information invisible and inaudible to all except the authorised parties who possess the key.
@@
Consider a human-centric audiovisual content in which an individual’s speech was recorded both aurally and visually. Suppose that the message is hidden within the pixel intensities or sound waves, serving an arbitrary purpose. Although steganographic systems can be designed to withstand distortion caused by common signal processing operations such as resampling, there remains a risk of unforeseen manipulations that go beyond the robustness premise. Generative AI technologies, such as face-swapping and voice-cloning, can potentially wipe out the hidden message, as the synthetic content has a non-trivial chance of overwriting the subtle changes made for the purpose of steganography.
@@
If the signals in both spatial and temporal domains are at risk of unforeseen manipulations, it prompts a reconsideration of what remains truly invariant. In quest for such an invariant domain, this study proposes a paradigm in steganography where messages are concealed beyond spatial and temporal dimensions. Motivated by multi-agent collaboration [61], we develop a chain of multimodal AI that deconstructs audiovisual content into cover text, embeds messages within the linguistic domain, and then reconstructs the audiovisual content by synchronising visual and auditory modalities with the stego text. In essence, our methodology transitions into the linguistic domain of audiovisual content, which is inherently invariant to direct spatiotemporal interference. The foundation of applied linguistic steganographic coding involves biasing the statistical distribution of word sampling [62]. Suppose a set of keywords (tokens) is shared between the sender and the receiver as the stego key. The message is encoded by biasing the word sampling process of a language generation model tasked with paraphrasing the given text, steering the model towards generating words from the shared set. The message can later be decoded by analysing the statistical likelihood of word choices, in comparison to the expected distribution of the shared set. The accuracy of message transmission is evaluated in both zero-bit and multi-bit capacity settings. In contrast to the perceptual similarity typically adopted in prior work, fidelity is extended to encompass both biometric and semantic similarities, representing identities of the recorded face and voice, as well as the ideas conveyed through the media. Secrecy is assessed through statistical analysis of cover and stego texts. Robustness is tested under conditions such as audiovisual resampling, face-swapping, voice-cloning, and combined scenarios.
@@
The remainder of this paper is organised as follows. Section Preliminaries outlines the scope of this study in relation to the primary properties of steganography. Section Methodology presents the proposed methodology, which involves a chain of multimodal AI. Section Evaluations details the experimental implementation and discusses the evaluation results. Finally, Section Conclusion concludes the paper with a summary of the research findings and potential directions for future research.
Preliminaries
@@
Fidelity
Fidelity refers to the degree of similarity between the cover and stego versions of the medium. A typical fidelity requirement is based on perceptual similarity, guaranteeing that any distortion caused by the steganographic system is either invisible or inaudible to human perception. The steganographic system should not compromise the integrity of the auditory or visual quality to the point of being noticeable to the listener or viewer. In a broader sense, however, the concept of fidelity can be relaxed to refer to semantic similarity, where the alterations made to the medium may be noticeable but do not change the fundamental meaning or purpose conveyed. This is often exemplified in linguistic steganography, where changes in lexicon, syntax, or even language may result in synonymous expressions or paraphrases that, while recognisable to a human observer, retain the core idea and intent of the original text. For human-centric audiovisual media, fidelity can be broadened to encompass biometric similarity, including face identity for visual content and voice identity for auditory content. This study focuses on both semantic and biometric similarities in the context of multimodal steganography.
Secrecy
@@
Robustness
Robustness refers to the ability to reliably decode the message even in the presence of unintentional degradations or malicious operations. It is pivotal in applications such as fact-checking and cyber-forensics, as it enables the tracing of a content’s origin and provides a means to verify whether the media has been fabricated or misrepresented. Central to achieving robustness is the principle of orthogonality, which ensures that the embedded message remains unaffected and detectable despite a range of potential manipulations. When the message is embedded in a domain that is either orthogonal or invariant to the potential attacks, it remains immune to their effects. For example, a visual or auditory medium can be projected into the frequency domain, where the message is embedded in parts of the frequency spectrum that are less likely to be compromised by potential manipulations. This study exploits the orthogonality between the linguistic and audiovisual domains, allowing the message to withstand changes in either the auditory or visual modality.
@@
Message Encoding Process
The encoding process begins with demultiplexing the cover multimedia container into cover video and cover audio. Next, the cover audio is transcribed into cover text, which is then encoded into stego text using the shared key. In the zero-bit capacity setting, only one key $K$ is shared. In the multi-bit capacity setting, the key $K$ to be used is selected from a shared key set based on the corresponding message symbol $S$, where the size of the message alphabet $\|\mathcal{S}\|$ matches the size of the key set $\|\mathcal{K}\|$. The stego text is then narrated into stego audio, and the cover video is synchronised with the stego audio. Finally, both stego video and stego audio are multiplexed into a stego multimedia container. An overview of the message encoding process is illustrated in Figure 1.
@@
The cover audio stream $A$ is transcribed into the cover text transcript $T$ using a speech-to-text (STT) model:
@@
Consider a language generation model prompted to paraphrase the given text. The word sampling process during paraphrasing is parameterised by a shared key $K$, selected based on the intended message symbol. The language generation model paraphrases the cover text transcript $T$ and results in a stego text transcript $T^{\prime}$:
@@
The word sampling process is biased toward generating tokens in the keyword set $\mathcal{W}$. The output of a language generation model is a sequence of logits, representing unnormalised probabilities of selecting each token from the dictionary. Let $z_{w}$ denote the logit for the token $w$. The probability of selection is biased by adjusting the logit:
@@
$p_{w}=\frac{\exp(z^{\prime}_{w})}{\sum_{v\in\mathcal{V}}\exp(z^{\prime}_{v})}.$ (6)
As long as the size of the keyword set is large enough, it is possible to preserve semantic equivalence between the cover and stego transcripts, with a statistically significant portion of contextual synonyms sampled from the keyword set.
@@
The stego text transcript $T^{\prime}$ is converted into the auditory modality using the voice cloned from the cover audio stream $A$ through a text-to-speech (TTS) model:
@@
To maintain audiovisual consistency, the stego audio stream $A^{\prime}$ is synchronised with the cover video stream $V$ using a lip-synchronisation model:
@@
Message Decoding Process
@@
The query audio stream $\hat{A}$ is transcribed into a query text transcript $\hat{T}$ using an STT model:
@@
Figure 3: Evaluation of accuracy in zero-bit capacity setting.
@@
Figure 4: Evaluation of accuracy in multi-bit capacity setting.
Evaluations
@@
  • Language Generator: An open-source language generation model named Llama developed by Meta [66].
These models were integrated to enable a fully automated steganographic system capable of generating semantics-aligned identity-preserved audiovisual media. The applied LLama model is version 3.2, with instruction-tuning and 1 billion parameters. At the time of writing, it represents a state-of-the-art release with a minimal model size, making it suitable for deployment and execution on lightweight hardware for reasoning tasks. The temperature of language generation was set to 0.1 to concentrate the sampling process on the most probable tokens, resulting in less variety in responses but higher coherence. The evaluations on zero-bit capacity, fidelity and robustness were conducted on a 9-second speech by C. E. Shannon from the MIT centennial film ‘The Thinking Machine’ [67], which served as the cover medium. The evaluations on multi-bit capacity and secrecy were conducted on 100 paragraphs of approximately 50 words each from ‘Alice’s Adventures in Wonderland’ by Lewis Carroll.
Evaluation of Zero-Bit Capacity
@@
Evaluation of Multi-Bit Capacity
@@
Figure 6: Evaluation of secrecy with respect to perplexity.
Evaluation of Fidelity
@@
Evaluation of Secrecy
At first glance, the challenge of steganalysis may appear to overlap with generative AI detection, as the proposed system involves a chain of AI to generate stego objects. However, as generative AI becomes increasingly prevalent in society, AI-synthesised content alone does not necessarily imply concealed communication. Consequently, classifying all AI-generated content as stego objects would be impractical. Due to the lack of existing steganalysis tools tailored to our proposed system, we evaluate the statistical deviations between cover and stego texts. Figure 7 shows the frequency distribution of words in the cover and stego texts, plotted on a log-log scale. The distribution for both cover and stego texts follows a similar pattern, with a steep decline in frequency as rank increases, typical of a Zipfian distribution [71]. Figure 6 presents the perplexity distribution for both cover and stego texts, measured using the Llama language model. Both cover and stego texts exhibit a similar range and central tendency in perplexity values, although the stego texts contain more outliers. These findings suggest that, as long as the stego texts are not outliers, they can be sampled with the same level of uncertainty as the cover texts, as reflected in their perplexity distribution. The results further support the statistical indistinguishability of the stego content.
@@
Evaluation of Robustness
Figure 8 compares the robustness of the proposed steganographic system with the visual steganographic system HiDDeN [72] and the auditory steganographic system WavMark [73] in terms of message extraction accuracy (complementary to BER). The evaluation was conducted under four conditions: no manipulation, resampling, deepfake, and hybrid, which combines both resampling and deepfake manipulations. The resampling algorithm reduced the display resolution to 25%, from 320 × 240 to 80×60 pixels, halved the video frame rate from 14 to 7 frames per second (fps), and halved the audio sample rate from 24 to 12 kilohertz (kHz). Deepfake video streams were generated using the face-swapping model SimSwap [74], with 100 faces from the FaceForensics++ dataset [75]. Deepfake audio streams were generated using the voice-cloning model XTTS [64], with 100 voices from the LibriSpeech corpus [76]. Our system was configured to operate at a capacity of one bit per audiovisual content. When no modification was applied, all three systems achieved perfect accuracy, confirming their ability to reliably communicate messages in an ideal setting. However, when subjected to resampling, HiDDeN exhibited a slight decline in accuracy, dropping to 0.90, whereas WavMark and our system remained unaffected. Under deepfake transformations, both HiDDeN and WavMark experienced further performance degradation, with accuracy dropping to 0.73 and 0.71, respectively, while our method maintained perfect robustness. The hybrid scenario, which introduced both resampling and deepfake manipulations, resulted in the lowest accuracy for HiDDeN at 0.71 and WavMark at 0.68, yet our system remained entirely resilient. Overall, these results demonstrate the vulnerability of existing systems to various forms of distortion, whereas our method consistently demonstrates superior robustness across all tested scenarios. Figure 9 visualises a pair of cover and stego media, along with the resampling and deepfake variants. It can be observed that while the cover transcript was altered by the steganographic system, the core semantics were preserved in the stego transcript. For instance, the text changed from ‘I confidently expect that within a matter of 10 or 15 years, something will emerge from the laboratories, which is not too far from the robot of science fiction fame’ to ‘Scientists are likely to create a new artificial life form that closely resembles the intelligent machines of science fiction’.
@@
Figure 9: Visualisation of cover, stego, resampling and deepfake versions of media across visual, auditory and linguistic modalities.
Conclusion
In conclusion, this study investigates the vulnerability of steganography to the evolving capabilities of generative AI, with a focus on the risk of synthetic content overwriting hidden messages. Instead of relying on the spatial or temporal domains, messages are embedded in the linguistic domain of audiovisual content based on a chain of multimodal AI. At the core of this chain is a language generation model tasked with paraphrasing the given transcript, while the word sampling process is biased towards a shared set of tokens. The paraphrased transcript is then converted into sound using a voice-cloning model, and the video is aligned with the audio using a lip-synchronisation model. Several aspects of the proposed steganographic system are evaluated, including accuracy, fidelity, secrecy and robustness, across various metrics and conditions.
The proposed system is inherently composed of a tightly coupled chain of AI components, each dependent on the output of the previous stage. While this sequential collaboration enables coherent synthesis of audiovisual stego media, it also introduces a high degree of interdependence, which may affect system flexibility and fault tolerance. A failure or degradation in an antecedent module can propagate downstream, potentially compromising stability in both encoding and decoding. This architectural choice prioritises semantic alignment but may not represent the optimal strategy in all contexts. An alternative design could involve parallel multimodal embeddings to enhance robustness and enable message extraction and verification across different modalities. In other words, future research may explore the integration of steganographic methods for individual modalities into a unified framework, where they complement each other through the cooperation and interaction of multimodal AI systems.